# Compact manifolds of dimension $n\geq 12$ with positive isotropic curvature

### v21

@@ 2 Orbifold ancient $\kappa$-solutions and standard solutions
Proof.   The arguments are adapted from the proof of Theorem 6.18 in [20], Theorems 3.9 in [29], and Proposition 3.6 in [51]. Note that by Corollary 6.7 in [20] (see also the first paragraph in the proof of Proposition 2.5 here) we know that $(\mathcal{O},g(t))$ is strictly PIC2. By an orbifold version of Gromoll-Meyer’s soul theorem (see Proposition 3.4 in [51]), $\mathcal{O}$ is diffeomorphic to the quotient orbifold $\mathbb{R}^{n}//\Gamma$ for some finite subgroup $\Gamma$ of $O(n)$ acting freely on $\mathbb{S}^{n-1}$. So we have a $\Gamma$-invariant projection $\pi:\mathbb{R}^{n}\rightarrow\mathcal{O}$. We pull back the Ricci flow $(\mathcal{O},g(t))$ to $(\mathbb{R}^{n},\tilde{g}(t))$ via $\pi$, which is then a $\Gamma$-invariant ancient $\kappa$-solution on a smooth manifold. Here the $\kappa$-noncollapsing property of $(\mathbb{R}^{n},\tilde{g}(t))$ follows from the $\kappa$-noncollapsing assumption for $(\mathcal{O},g(t))$ and the fact that $\pi|_{B(\tilde{p},t,r)}:B(\tilde{p},t,r)\rightarrow B(p,t,r)$ is surjective for any $p\in\mathcal{O}$, $\tilde{p}\in\pi^{-1}(p)$, any $t$ (for which $g(t)$ is defined), and any $r>0$ (see for example the proof of Proposition 3.7 in [29]). We fix $t_{0}=0$. For any small $\eta>0$, let $M_{\eta}$ be the set of all points in $\mathbb{R}^{n}$ which are not the centers of $\eta$-necks in $(\mathbb{R}^{n},\tilde{g}(0))$. From Step 1 in the proof of Theorem 6.18 in [20] we know that $M_{\eta}$ is compact. Let $y$ be any point in $\partial M_{\eta}$. From Step 2 in the proof of Theorem 6.18 in [20] we have
@@
for all $x\in\mathbb{R}^{n}$ and $t\leq 0$, where $\omega:[0,\infty)\rightarrow[0,\infty)$ is a function independent of $\kappa$. (For the proof, instead of Theorem 6.12 in [20] one can also use an extension of Corollary 9.62 in [67] to our situation, which follows from the fact that the asymptotic volume ratio of any ancient $\kappa$-solution (see Definition 6.10 in [20]) is also zero in the higher dimensional case. This fact can be shown by using Theorem 6.1 and Proposition 6.5 in [20].) Note that $y$ is the center of a $2\eta$-neck in $(\mathbb{R}^{n},\tilde{g}(0))$, and by [48] this neck (away from the two ends) is foliated by almost horizontal constant mean curvature spheres (of dimension $n-1$) in a unique way. Let $\tilde{\Sigma}_{y}$ be the leaf of this foliation which passes through $y$. By Theorem G1.1 in [48] (whose proof applies to any dimension $n\geq 4$) (of course, we can also use the solution of Schoenflies conjecture in dimensions $n\geq 5$ here), $\tilde{\Sigma}_{y}$ bounds a compact domain $\tilde{\Omega}_{y}$ (in $\mathbb{R}^{n}$) diffeomorphic to a standard disk $D^{n}$. By Step 3 in the proof of Theorem 6.18 in [20] $R(y,0)\hskip 1.13809pt\text{diam}_{\tilde{g}(0)}(\tilde{\Omega}_{y})^{2}\leq C$, where $C=C(n,\theta,\eta)$ is independent of $\kappa$.
@@
As in the proof of Proposition 3.4 in [31], there is a point $y\in\partial M_{\eta}$ such that $M_{\eta}\subset\tilde{\Omega}_{y}$. Fix such a point $y$. We rescale the solution $\tilde{g}(t)$ so that $R(y,0)=1$ after rescaling; the rescaled solution will still be denoted by $\tilde{g}(t)$. Note that $\mathbb{R}^{n}$ ($n\geq 5$) has only one end. For $\eta>0$ sufficiently small (depending on $\varepsilon$), we can use Lemma 5.6 and Proposition 5.7 (see also Lemma C2.1, Corollary C2.3 and Theorem C2.4 in [48]) repeatedly to get a Hamilton’s canonical parametrization $\tilde{\phi}:\mathbb{S}^{n-1}\times(a,+\infty)\rightarrow\widetilde{\mathcal{T}}\subset\mathbb{R}^{n}$ which gives an $\varepsilon$-tube $\widetilde{\mathcal{T}}$ in $(\mathbb{R}^{n},\tilde{g}(0))$ containing $\mathbb{R}^{n}\setminus B(y,0,\sqrt{C}+1)$, where $B(y,0,\sqrt{C}+1)$ denotes the ball of radius $\sqrt{C}+1$ w.r.t. the rescaled pull-back metric at time 0 centered at $y$. We will fix such an $\eta$. We take the Hamilton’s canonical parametrization $\tilde{\phi}$ to be maximal in the sense that $\widetilde{\mathcal{T}}$ cannot be properly included in the image of any Hamilton’s canonical parametrization which gives an $\varepsilon$-tube in $(\mathbb{R}^{n},\tilde{g}(0))$.
@@
As in the proof of Lemma 5.6, the parametrization $\tilde{\phi}$ above can be pushed down to get a Hamilton’s canonical parametrization $\phi:\mathbb{S}^{n-1}/\Gamma\times(a,+\infty)\rightarrow\mathcal{T}\subset\mathcal{O}$ which gives an $\varepsilon$-tube in $(\mathcal{O},g(0))$. (Compare the proof of Theorem 3.9 in [29]; here our proof is somewhat different.)
By Proposition 4.3 in [41] (see also p. 967-968 in [51]), there exists a point in $\mathbb{R}^{n}$ which is a soul of $(\mathbb{R}^{n},\tilde{g}(0))$ and which is fixed by $\Gamma$. We denote this point by $O$. As in [51] one can show that $O$ has distance $\leq\sqrt{C}+1$ from $y$ at time 0.
Let $S\in\mathcal{O}$ be the image of $O$ under the projection $\pi$. For any point $x\in\mathcal{O}$ with $d_{0}(x,S)\geq 2\sqrt{C}+1$, we can use the $\phi$ above to give an $\varepsilon$-neck centered at $x$.
Then we want to find an $\varepsilon$-cap diffeomorphic to $\mathbb{R}^{n}//\Gamma$ in $\mathcal{O}$. Let $\tilde{z}_{0}\in\mathbb{R}^{n}$ be a point with $d_{0}(O,\tilde{z}_{0})=10\sqrt{C}$. Since $d_{0}(O,y)\leq\sqrt{C}+1$, we see that $8\sqrt{C}<d_{0}(\tilde{z}_{0},y)<12\sqrt{C}$. We denote the almost horizontal constant mean curvature $(n-1)$-sphere passing through $\tilde{z}_{0}$ by $\tilde{\Sigma}$, which bounds a compact domain $\tilde{\Omega}$ (in $\mathbb{R}^{n}$) diffeomorphic to a closed, standard $n$-disk $D^{n}$. From above we know that $\tilde{\Omega}$ is $\Gamma$-invariant, and $\text{Int}\hskip 1.42262pt\tilde{\Omega}//\Gamma$ contains an $\varepsilon$-neck $W$ near its end. As in the proof of Proposition 3.6 in [51] we can use the heat equation method (cf. Appendixes F and G in [48] and the proof of Theorem 3.9 in [29]) to smooth the Busemann function based at the point $O$, and show that $\tilde{\Omega}$ is $\Gamma$-equivariantly diffeomorphic to $D^{n}$; cf. also the proof of Proposition 5.4 in the appendix. (By the way, we would like to point out that the reaction term on the RHS of the evolution equation for the Hessian of $u$ (where $u$ is a solution to the heat equation coupled with the Ricci flow) on p. 56 of [29] satisfies the so called null-eigenvector assumption (Definition 10.6 in [35]) under the PIC2 condition.) So $\text{Int}\hskip 1.42262pt\tilde{\Omega}//\Gamma$ is diffeomorphic to $\mathbb{R}^{n}//\Gamma$, and is an $\varepsilon$-cap; in the case that $\Gamma$ is trivial, that $\text{Int}\hskip 1.42262pt\tilde{\Omega}$ has the additional property in the definition of $\varepsilon$-cap is guaranteed by Proposition 2.6. From above, the Harnack inequality (Corollary 6.3 in [20]), and the fact that $\pi|_{B(\tilde{p},0,r)}:B(\tilde{p},0,r)\rightarrow B(p,0,r)$ is surjective for any $p\in\mathcal{O}$, $\tilde{p}\in\pi^{-1}(p)$ and any $r>0$, we see that $C^{\prime-1}R(y,0)\leq R(z,0)\leq C^{\prime}R(y,0)$ and $\text{diam}_{g(0)}(\tilde{\Omega}//\Gamma)\leq C^{\prime}R(z,0)^{-\frac{1}{2}}$ for any $z\in\tilde{\Omega}//\Gamma$, where $C^{\prime}=C^{\prime}(n,\theta,\varepsilon)$ is independent of $\kappa$.
@@
Let $V:=(\text{Int}\hskip 1.42262pt\tilde{\Omega}//\Gamma)\setminus W$. If $x_{0}\in\text{Int}\hskip 1.42262ptV$, $\text{Int}\hskip 1.42262pt\tilde{\Omega}//\Gamma$ is an $\varepsilon$-cap centered at $x_{0}$. If $x_{0}\in\mathcal{O}\setminus\text{Int}\hskip 1.42262ptV$, we can use a suitable restriction of $\phi$ to get an $\varepsilon$-neck centered at $x_{0}$. Now the existence of the desired constants $C_{1}$ and $C_{2}$ is clear. $\Box$
@@
There exist a positive function $\omega:[0,\infty)\rightarrow(0,\infty)$ and a positive constant $C_{3}$ both depending only on $n$ and $\theta$ with the following property. Assume that $(\mathcal{O},g(t)_{t\in(-\infty,0]})$ is an orbifold ancient $\kappa$-solution of dimension $n\geq 5$ (with at most isolated singularities) which satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in PIC$. Then
@@
Proof.   For (i), we use Proposition 3.4 in [51], an orbifold version of the sphere theorem in [23] (see the proof of Proposition 2.5 above) and Theorems 6.13 (in the proof of this theorem, instead of Theorem 6.12 in [20] one can also use an extension of Corollary 9.62 in [67] to our situation) and 6.19 in [20], and argue as in the proof of Proposition 3.7 in [29]. The conclusion in (ii) follows from (i) and Shi’s interior derivative estimates. $\Box$
@@
(cf. [72], Corollaries 6.20 and 6.22 in [20], and Theorem 3.10 in [29])   Given a small positive constant $\varepsilon$ and a constant $\theta>0$, there exist positive constants $C_{1}=C_{1}(n,\theta,\varepsilon)$, $C_{2}=C_{2}(n,\theta,\varepsilon)$ and $C_{3}=C_{3}(n,\theta)$ with the following property: Suppose $(\mathcal{O},g(t))$ is an orbifold ancient $\kappa$-solution of dimension $n\geq 5$ (with at most isolated singularities) which satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in PIC$. Then either $(\mathcal{O},g(t))$ is compact and strictly PIC2 for any $t$ (hence diffeomorphic to a spherical orbifold), or for each space-time point $(x_{0},t_{0})$, there is an open subset $U$ of $\mathcal{O}$ with $B(x_{0},t_{0},C_{1}^{-1}R(x_{0},t_{0})^{-\frac{1}{2}})\subset U\subset B(x_{0},t_{0},C_{1}R(x_{0},t_{0})^{-\frac{1}{2}})$ and with the values of the scalar curvature in $U$ at time $t_{0}$ lying (strictly) between $C_{2}^{-1}R(x_{0},t_{0})$ and $C_{2}R(x_{0},t_{0})$, which falls into one of the following two categories:
(a) $U$ is a strong $\varepsilon$-neck centered at $(x_{0},t_{0})$,
(b) $U$ is an $\varepsilon$-cap centered at $(x_{0},t_{0})$;
@@
Proof.   Let $(\mathcal{O},g(t))$ be as in the assumption. If $\mathcal{O}$ is compact, by the $\kappa$-noncollapsing assumption on all scales it cannot be isometric to a metric quotient of the round cylinder. Then by an orbifold version of Corollary 6.7 in [20] in this case $(\mathcal{O},g(t))$ is strictly PIC2 for any $t$, so $\mathcal{O}$ is diffeomorphic to a spherical orbifold. If $\mathcal{O}$ is noncompact, the result except for the derivative estimates follows from Propositions 2.5 and 2.7 and Lemma 5.8, and the derivative estimates follows from Proposition 2.8 (ii). $\Box$
@@
It turns out that orbifold standard solutions have canonical neighborhood property similar to that of orbifold ancient $\kappa$-solutions, except that an $\varepsilon$-neck in an orbifold standard solution may not be strong.

### v22

@@ 2 Orbifold ancient $\kappa$-solutions and standard solutions
Proof.   The arguments are adapted from the proof of Theorem 6.18 in [21], Theorems 3.9 in [31], and Proposition 3.6 in [54]. Note that by Corollary 6.7 in [21] (see also the first paragraph in the proof of Proposition 2.5 here) we know that $(\mathcal{O},g(t))$ is strictly PIC2. By an orbifold version of Gromoll-Meyer’s soul theorem (see Proposition 3.4 in [54]), $\mathcal{O}$ is diffeomorphic to the quotient orbifold $\mathbb{R}^{n}//\Gamma$ for some finite subgroup $\Gamma$ of $O(n)$ acting freely on $\mathbb{S}^{n-1}$. So we have a $\Gamma$-invariant projection $\pi:\mathbb{R}^{n}\rightarrow\mathcal{O}$. We pull back the Ricci flow $(\mathcal{O},g(t))$ to $(\mathbb{R}^{n},\tilde{g}(t))$ via $\pi$, which is then a $\Gamma$-invariant ancient $\kappa$-solution on a smooth manifold. Here the $\kappa$-noncollapsing property of $(\mathbb{R}^{n},\tilde{g}(t))$ follows from the $\kappa$-noncollapsing assumption for $(\mathcal{O},g(t))$ and the fact that $\pi|_{B(\tilde{p},t,r)}:B(\tilde{p},t,r)\rightarrow B(p,t,r)$ is surjective for any $p\in\mathcal{O}$, $\tilde{p}\in\pi^{-1}(p)$, any $t$ (for which $g(t)$ is defined), and any $r>0$ (see for example the proof of Proposition 3.7 in [31]). We fix $t_{0}=0$. For any small $\eta>0$, let $M_{\eta}$ be the set of all points in $\mathbb{R}^{n}$ which are not the centers of $\eta$-necks in $(\mathbb{R}^{n},\tilde{g}(0))$. Note that for $\eta$ sufficiently small, $M_{\eta}\neq\emptyset$ since it contains a soul of $(\mathbb{R}^{n},\tilde{g}(0))$; cf. for example p. 969 in [54] (or by using an extension of Lemma A.10 in [70]). From Step 1 in the proof of Theorem 6.18 in [21] we know that $M_{\eta}$ is compact. Let $y$ be any point in $\partial M_{\eta}$. From Step 2 in the proof of Theorem 6.18 in [21] we have
@@
for all $x\in\mathbb{R}^{n}$ and $t\leq 0$, where $\omega:[0,\infty)\rightarrow[0,\infty)$ is a function independent of $\kappa$. (For the proof, instead of using Theorem 6.12 in [21] (compare Remark 45.4 in [58]) one can also use an extension of Corollary 9.62 in [70] in our situation, which follows from the fact that the asymptotic volume ratio of any ancient $\kappa$-solution (see Definition 6.10 in [21]) is also zero in the higher dimensional case. This fact can be shown by using Theorem 6.1 and Proposition 6.5 in [21].) Note that $y$ is the center of a $2\eta$-neck in $(\mathbb{R}^{n},\tilde{g}(0))$, and by [51] this neck (away from the two ends) is foliated by almost horizontal constant mean curvature spheres (of dimension $n-1$) in a unique way. Let $\tilde{\Sigma}_{y}$ be the leaf of this foliation which passes through $y$. By Theorem G1.1 in [51] (whose proof applies to any dimension $n\geq 4$) (of course, we can also use the solution of Schoenflies conjecture in dimensions $n\geq 5$ here), $\tilde{\Sigma}_{y}$ bounds a compact domain $\tilde{\Omega}_{y}$ (in $\mathbb{R}^{n}$) diffeomorphic to a standard disk $D^{n}$. By Step 3 in the proof of Theorem 6.18 in [21] $R(y,0)\hskip 1.13809pt\text{diam}_{\tilde{g}(0)}(\tilde{\Omega}_{y})^{2}\leq C$, where $C=C(n,\theta,\eta)$ is independent of $\kappa$.
@@
As in the proof of Proposition 3.4 in [33], there is a point $y\in\partial M_{\eta}$ such that $M_{\eta}\subset\tilde{\Omega}_{y}$. Fix such a point $y$. We rescale the solution $\tilde{g}(t)$ so that $R(y,0)=1$ after rescaling; the rescaled solution will still be denoted by $\tilde{g}(t)$. Note that $\mathbb{R}^{n}$ ($n\geq 5$) has only one end. For $\eta>0$ sufficiently small (depending on $\varepsilon$), we can use Lemma A.6 and Proposition A.7 (see also Lemma C2.1, Corollary C2.3 and Theorem C2.4 in [51]) repeatedly to get a Hamilton’s canonical parametrization $\tilde{\phi}:\mathbb{S}^{n-1}\times(a,+\infty)\rightarrow\widetilde{\mathcal{T}}\subset\mathbb{R}^{n}$ which gives an $\varepsilon$-tube $\widetilde{\mathcal{T}}$ in $(\mathbb{R}^{n},\tilde{g}(0))$ containing $\mathbb{R}^{n}\setminus\text{Int}\hskip 1.42262pt\tilde{\Omega}_{y}$ such that $y$ is a center of an $\varepsilon$-neck (with Hamilton’s canonical parametrization) contained in $\widetilde{\mathcal{T}}$. Recall that by Proposition 4.3 in [44] (see also p. 967-968 in [54]), there exists a point in $\mathbb{R}^{n}$ which is a soul of $(\mathbb{R}^{n},\tilde{g}(0))$ and which is fixed by $\Gamma$. We denote this point by $O$. As in p. 969 of [54] one can show that $O$ is not in the nine-tenths of any $2\eta$-neck in $(\mathbb{R}^{n},\tilde{g}(0))$ if $\eta$ is sufficiently small. In particular, $O\in\text{Int}\hskip 1.42262ptM_{\eta}\subset\text{Int}\hskip 1.42262pt\tilde{\Omega}_{y}$. We will fix an $\eta$ satisfying the desired properties above. We take the Hamilton’s canonical parametrization $\tilde{\phi}$ to be maximal in the sense that $\widetilde{\mathcal{T}}$ cannot be properly included in the image of any Hamilton’s canonical parametrization which gives an $\varepsilon$-tube in $(\mathbb{R}^{n},\tilde{g}(0))$.
@@
As in the proof of Lemma A.6, the parametrization $\tilde{\phi}$ above can be pushed down to get a Hamilton’s canonical parametrization $\phi:\mathbb{S}^{n-1}/\Gamma\times(a,+\infty)\rightarrow\mathcal{T}\subset\mathcal{O}$ which gives an $\varepsilon$-tube in $(\mathcal{O},g(0))$. (Compare the proof of Theorem 3.9 in [31]; here our proof is somewhat different.)
@@
For any point $x\in\mathcal{O}\setminus\pi(\text{Int}\hskip 1.42262pt\tilde{\Omega}_{y})$, we can use the $\phi$ above to give an $\varepsilon$-neck centered at $x$.
For any $z\in\mathbb{R}^{n}\setminus\text{Int}\hskip 1.42262pt\tilde{\Omega}_{y}$ we denote the almost horizontal constant mean curvature $(n-1)$-sphere passing through $z$ by $\tilde{\Sigma}_{z}$, which bounds a compact domain $\tilde{\Omega}_{z}$ (in $\mathbb{R}^{n}$) diffeomorphic to a closed, standard $n$-disk $D^{n}$. Let $\tilde{W}$ be the $\varepsilon$-neck with Hamilton’s canonical parametrization which is contained in $\mathbb{R}^{n}\setminus\tilde{\Omega}_{y}$ such that $\tilde{\Sigma}_{y}$ is one connected component of the boundary of the closure of $\tilde{W}$. Let $z_{0}$ be any point in the connected component of the boundary of the closure of $\tilde{W}$ which is different from $\tilde{\Sigma}_{y}$. From above we know that $\tilde{\Omega}_{z_{0}}$ is $\Gamma$-invariant, and $\text{Int}\hskip 1.42262pt\tilde{\Omega}_{z_{0}}//\Gamma$ contains the $\varepsilon$-neck $W:=\pi(\tilde{W})$ near its end with $\partial\overline{W}\cap\text{Int}\hskip 1.42262pt\tilde{\Omega}_{z_{0}}//\Gamma$ ($=\pi(\tilde{\Sigma}_{y})$) a central cross-section of an $\varepsilon$-neck. Now we can use the heat equation method (cf. Appendixes F and G in [51] and the proof of Theorem 3.9 in [31]) to smooth the Busemann function based at the $\Gamma$-invariant soul $O$ a bit to a $\Gamma$-invariant smooth function $\tilde{b}$, and fuse $\tilde{b}$ with the $\Gamma$-invariant height function on the $\varepsilon$-neck centered at $z_{0}$ and parametrized by certain restriction of the map $\tilde{\phi}$ (or this height function multiplied by a constant and/or shifted by a constant if necessary) using a suitable cut-off function, and an equivariant Morse theoretical argument using (roughly speaking) the gradient flow of the resulting function to show that $\tilde{\Omega}_{z_{0}}$ is $\Gamma$-equivariantly diffeomorphic to $D^{n}$ (if $\varepsilon$ is small enough, which we may assume); compare the proof of Theorem G1.1 in [51], cf. also the proof of Theorem 3.9 in [31], Proposition 3.6 in [54] and Proposition A.4 in the appendix. (By the way, we would like to point out that the reaction term on the RHS of the evolution equation for the Hessian of $u$ (where $u$ is a solution to the heat equation coupled with the Ricci flow) on p. 56 of [31] satisfies the so called null-eigenvector assumption (Definition 10.6 in [38]) under the PIC2 condition.) So $\text{Int}\hskip 1.42262pt\tilde{\Omega}_{z_{0}}//\Gamma$ is diffeomorphic to $\mathbb{R}^{n}//\Gamma$, and is an $\varepsilon$-cap; in the case that $\Gamma$ is trivial, that $\text{Int}\hskip 1.42262pt\tilde{\Omega}_{z_{0}}$ has the additional property in the definition of $\varepsilon$-cap is guaranteed by Proposition 2.6.
From above, the Harnack inequality (Corollary 6.3 in [21]), and the fact that $\pi|_{B(\tilde{p},0,r)}:B(\tilde{p},0,r)\rightarrow B(p,0,r)$ is surjective for any $p\in\mathcal{O}$, $\tilde{p}\in\pi^{-1}(p)$ and any $r>0$, we see that $C^{\prime-1}R(y,0)\leq R(\bar{z},0)\leq C^{\prime}R(y,0)$ and $\text{diam}_{g(0)}(\tilde{\Omega}_{z_{0}}//\Gamma)\leq C^{\prime}R(\bar{z},0)^{-\frac{1}{2}}$ for any $\bar{z}\in\tilde{\Omega}_{z_{0}}//\Gamma$, where $C^{\prime}=C^{\prime}(n,\theta,\varepsilon)$ is a positive constant independent of $\kappa$. In particular, $R(\bar{z},0)$ and $R(\bar{z}^{\prime},0)$ are comparable if both $\bar{z}$ and $\bar{z}^{\prime}$ are points in $\tilde{\Omega}_{z_{0}}//\Gamma$. We also observe that if $\tilde{z}$ is a center of an $\varepsilon$-neck (for the metric $\tilde{g}(0)$) contained in the end of $\tilde{\Omega}_{z_{0}}$,
$\text{vol}_{\tilde{g}(0)}(\tilde{\Omega}_{z_{0}})>\text{vol}_{\tilde{g}(0)}(B_{\tilde{g}(0)}(\tilde{z},R(\tilde{z},0)^{-1/2}))\geq\mu(n)R(\tilde{z},0)^{-n/2},$
where $\mu(n)$ is a positive constant depending on $n$. If $x_{0}\in\pi(\text{Int}\hskip 1.42262pt(\tilde{\Omega}_{y}))$, $\text{Int}\hskip 1.42262pt\tilde{\Omega}_{z_{0}}//\Gamma$ is an $\varepsilon$-cap centered at $x_{0}$. Now the existence of the desired constants $C_{1}$ and $C_{2}$ is clear. $\Box$
@@
There exists a positive function $\omega:[0,\infty)\rightarrow(0,\infty)$ and a positive constant $C_{3}$ both depending only on $n$ and $\theta$ with the following property. Assume that $(\mathcal{O},g(t)_{t\in(-\infty,0]})$ is an orbifold ancient $\kappa$-solution of dimension $n\geq 5$ (with at most isolated singularities) which satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}$. Then
@@
Proof.   For (i), we use Proposition 3.4 in [54], an orbifold version of the sphere theorem in [24] (see the proof of Proposition 2.5 above) and Theorems 6.13 (in the proof of this theorem, instead of Theorem 6.12 in [21] one can also use an extension of Corollary 9.62 in [70] in our situation) and 6.19 in [21], and argue as in the proof of Proposition 3.7 in [31]. The conclusion in (ii) follows from (i) and Shi’s interior derivative estimates. $\Box$
@@
(cf. [76], Corollaries 6.20 and 6.22 in [21], and Theorem 3.10 in [31])   Given a small positive constant $\varepsilon$ and a constant $\theta>0$, there exist positive constants $C_{1}=C_{1}(n,\theta,\varepsilon)$, $C_{2}=C_{2}(n,\theta,\varepsilon)$ and $C_{3}=C_{3}(n,\theta)$ with the following property: Suppose $(\mathcal{O},g(t))$ is an orbifold ancient $\kappa$-solution of dimension $n\geq 5$ (with at most isolated singularities) which satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}$. Then either $(\mathcal{O},g(t))$ is compact and strictly PIC2 for any $t$ (hence diffeomorphic to a spherical orbifold), or for each space-time point $(x_{0},t_{0})$, there is an open subset $U$ of $\mathcal{O}$ with $\overline{B(x_{0},t_{0},C_{1}^{-1}R(x_{0},t_{0})^{-\frac{1}{2}})}\subset U\subset B(x_{0},t_{0},C_{1}R(x_{0},t_{0})^{-\frac{1}{2}})$ and with the values of the scalar curvature in $U$ at time $t_{0}$ lying (strictly) between $C_{2}^{-1}R(x_{0},t_{0})$ and $C_{2}R(x_{0},t_{0})$, which falls into one of the following two categories:
(a) $U$ is a strong $\varepsilon$-neck diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(-1,1)$ centered at $(x_{0},t_{0})$, or
(b) $U$ is an $\varepsilon$-cap centered at $(x_{0},t_{0})$ with any cross-section of the $\varepsilon$-neck contained in it diffeomorphic to $\mathbb{S}^{n-1}/\Gamma$
for some finite subgroup $\Gamma$ of $O(n)$ acting freely on $\mathbb{S}^{n-1}$, and
$vol_{g(t_{0})}(U)>(C_{2}|\Gamma|)^{-1}R(x_{0},t_{0})^{-n/2};$
@@
Proof.   Let $(\mathcal{O},g(t))$ be as in the assumption. If $\mathcal{O}$ is compact, by the $\kappa$-noncollapsing assumption on all scales it cannot be isometric to a metric quotient of the round cylinder. Then by an orbifold version of Corollary 6.7 in [21] in this case $(\mathcal{O},g(t))$ is strictly PIC2 for any $t$, so $\mathcal{O}$ is diffeomorphic to a spherical orbifold. If $\mathcal{O}$ is noncompact, the result except for the derivative estimates follows from Propositions 2.5 and 2.7 and Lemma A.8, and the derivative estimates follows from Proposition 2.8 (ii). $\Box$
@@
It turns out that orbifold standard solutions (see the appendix) have canonical neighborhood property similar to that of orbifold ancient $\kappa$-solutions, except that an $\varepsilon$-neck in an orbifold standard solution may not be strong.